# Riemann-Silberstein geometric phase for high-dimensional light manipulation

### v1

@@ Abstract
Geometric phases provide a powerful mechanism for light manipulation. In particular, the Pancharatnam-Berry (PB) phase has enabled optical metasurfaces with broad applications. However, the PB phase is based on polarization evolution in a two-dimensional space, which fails to account for other polarization degrees of freedom. Here, we generalize the concept of geometric phase to a four-dimensional (4D) Riemann-Silberstein (RS) space that characterizes the complete electromagnetic polarization, including electric, magnetic, and hybrid polarizations. We show that the 4D polarization evolution in the RS space can give rise to a new geometric phase—the RS phase—in addition to the PB phase. The PB phase depends on optical spin and usually manifests in circularly polarized light, whereas the RS phase depends on optical linear momentum and can manifest in arbitrarily polarized light. Their synergy provides a unified geometric framework for light propagation at interfaces and enables unprecedented high-dimensional light control. As a proof of principle, we propose and demonstrate RS metasurfaces capable of multiplexed wavefront shaping, which can reconfigure up to twelve distinct outputs via switching incident 4D polarization. Our work uncovers a new class of optical geometric phases, with promising applications in high-capacity optical communication, parallel information processing, and multifunctional nanophotonic design.
@@ Introduction
Geometric phases emerge from state evolution in parameter space and provide a unified framework for understanding fundamental phenomena in quantum and classical physics [1, 2, 3, 4]. These phases have been extensively studied in diverse physical systems, including quantum particles [2, 5, 6, 7], condensed matter [8, 9], and classical wave systems [10, 11, 12, 13]. In optics, geometric phases can give rise to intriguing phenomena such as spin-orbit interactions [14] and photonic topological states [15, 16, 17, 18], providing profound insights into the geometric and topological properties of optical fields [19, 20, 21, 22] and enabling novel mechanisms for light manipulation [23, 24, 25, 26, 27, 28].
The optical PB phase has recently attracted significant attention for its crucial role in wavefront manipulation by metasurfaces [29, 30, 31, 32, 33, 34, 35, 36]. This phase arises from SU(2) polarization evolution on the Poincaré sphere [37, 38], which is a 2D space describing the polarization of a two-component vector field. However, light is an electromagnetic wave comprising an electric field $\mathbf{E}$ and a magnetic field $\mathbf{H}$, which form a bispinor in the Dirac-like framework [39, 40]. For monochromatic waves, both $\mathbf{E}$ and $\mathbf{H}$ are two-component vector fields within the local frame defined by their polarization ellipses [19, 41], and their polarizations can be different. Consequently, the complete polarization state of a monochromatic electromagnetic wave resides in a 4D Hilbert space (i.e., direct sum of electric and magnetic polarization spaces), termed RS space in this paper, which simultaneously characterizes electric, magnetic, and hybrid electric-magnetic polarizations. The SU(4) evolution of the 4D complete polarization can generate nontrivial geometric phases beyond the conventional PB phase, which has thus far remained elusive. These geometric phases are essential to understanding the complete vector properties of light and making full use of its degrees of freedom.
In this work, we extend the polarization space from 2D to 4D and investigate the geometric phase emerging from SU(4) evolution of the complete electromagnetic polarization. We identify a new class of geometric phase—the RS phase, which is governed by the hybrid electric-magnetic polarization, in contrast to the conventional PB phase associated with the individual electric (magnetic) polarization. The complementary nature of the RS and PB phases enables a geometric description of the phase shifts induced by light transmission and reflection at interfaces. Leveraging this dual-phase mechanism, we propose and demonstrate RS metasurfaces that can achieve high-dimensional wavefront manipulation, generating up to twelve distinct output wavefronts under the normal incidence of a plane wave with switchable 4D polarization.
@@
Results
4D polarization with Poincaré hypersphere representation
@@ Results
A monochromatic electromagnetic field in three-dimensional space can be described by the wave field $\bm{\Psi}=\mathbf{E}+\mathrm{i}\mathbf{H}$, where $\mathbf{E}(\mathbf{r},t)=E_{i}(\mathbf{r},t)\hat{\mathbf{e}}_{i}+E_{j}(\mathbf{r},t)\hat{\mathbf{e}}_{j}$ is the complex electric field in the local frame with bases $\left(\hat{\mathbf{e}}_{i},\hat{\mathbf{e}}_{j}\right)$, and $\mathbf{H}(\mathbf{r},t)=H_{i}(\mathbf{r},t)\hat{\mathbf{h}}_{i}+H_{j}(\mathbf{r},t)\hat{\mathbf{h}}_{j}$ is the complex magnetic field in the local frame with bases $(\hat{\mathbf{h}}_{i},\hat{\mathbf{h}}_{j})$. Gaussian units are used in the formulations. We note that $\mathbf{E}$ and $\mathbf{H}$ are in orthogonal spaces, i.e., they do not mix in the wave field $\bm{\Psi}$. The wave field $\bm{\Psi}$ exhibits both internal polarizations (i.e., the polarizations of individual $\mathbf{E}$ and $\mathbf{H}$ fields) and external polarization (i.e., the hybrid $\mathbf{E}$-$\mathbf{H}$ polarization), characterized by the 4D polarization state $|\Psi\rangle$ residing in the 4D RS space. This 4D polarization can be described by a Poincaré hypersphere [42, 43], as illustrated in Fig. 1a, which comprises three nested Poincaré spheres: E-sphere characterizing the $\mathbf{E}$ field polarization, H-sphere characterizing the $\mathbf{H}$ field polarization, and RS-sphere characterizing the hybrid $\mathbf{E}$-$\mathbf{H}$ polarization. The E-sphere has the north pole state $\left|N_{\mathrm{e}}\right\rangle$ corresponding to right-handed circularly-polarized (RCP) electric field and the south pole state $\left|S_{\mathrm{e}}\right\rangle$ corresponding to left-handed circularly-polarized (LCP) electric field; The H-sphere has the north pole state $\left|N_{\mathrm{m}}\right\rangle$ corresponding to RCP magnetic field and the south pole state $\left|S_{\mathrm{m}}\right\rangle$ corresponding to the LCP magnetic field. Any other point on the E-sphere or H-sphere denotes a $\mathrm{SU}(2)$ superposition state parameterized by the polar angle $\theta_{\mathrm{e},\mathrm{m}}$ and azimuthal angle $\phi_{\mathrm{e},\mathrm{m}}$. Notably, the RS-sphere characterizes the relationship between E-sphere and H-sphere with $\theta_{\mathrm{rs}}$ and $\phi_{\mathrm{rs}}$; its horizontal basis state $\left|H_{\mathrm{rs}}\right\rangle$ corresponds to an arbitrary state on the E-sphere; its vertical basis state $\left|V_{\mathrm{rs}}\right\rangle$ corresponds to an arbitrary state on the H-sphere. Any other point on the RS-sphere denotes a $\operatorname{SU}(4)$ superposition state $|\Psi\rangle$ (see Methods).
@@
In the following, we will focus on paraxial waves involving transverse electric and magnetic fields that are mutually perpendicular. In such cases, $\mathbf{E}$ and $\mathbf{H}$ exhibit the same polarization, and $\bm{\Psi}$ can be decomposed into a pair of orthogonal RS vectors: $\bm{\Psi}=\mathbf{F}_{1}+\mathbf{F}_{2}$, where $\mathbf{F}_{1}(\mathbf{r},t)=E_{i}(\mathbf{r},t)\hat{\mathbf{e}}_{i}+\mathrm{i}H_{j}(\mathbf{r},t)\hat{\mathbf{h}}_{j}$ and $\mathbf{F}_{2}(\mathbf{r},t)=E_{j}(\mathbf{r},t)\hat{\mathbf{e}}_{j}+\mathrm{i}H_{i}(\mathbf{r},t)\hat{\mathbf{h}}_{i}$. Consequently, the hybrid $\mathbf{E}$-$\mathbf{H}$ polarization reduces to the polarizations of the RS fields $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$. We note that the RS vectors $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ are composed of complex electric and magnetic fields, which are different from the conventional RS vector comprising real electric and magnetic fields [44, 45, 46, 47, 48].
@@
We use electromagnetic plane waves as an example to illustrate the 4D polarization and its representation on the Poincaré hypersphere. We first consider a plane wave with linearly polarized electric and magnetic fields: $\bm{\Psi}=(\hat{\mathbf{e}}_{x}+\mathrm{i}\kappa\hat{\mathbf{h}}_{y})Ee^{\mathrm{i}\kappa k_{0}z-\mathrm{i}\omega t}$, where $\kappa=\pm 1$ denotes the sign of the propagation direction relative to +z+z axis. Figures 1c and 1d show the instantaneous $\mathbf{E}$, $\mathbf{H}$, and $\mathbf{F}_{1}$ fields for $\kappa=1$ and $\kappa=-1$, respectively. The polarizations of $\mathbf{E}$ and $\mathbf{H}$ fields are represented by points “c” and “d” on the E-sphere and H-sphere in Fig. 1a. The RS field $\mathbf{F}_{1}$ is RCP for $\kappa=1$ and LCP for $\kappa=-1$, which are represented by points “c” and “d” on the RS-sphere in Fig. 1a. Naturally, we can introduce an RS spin density $\mathbf{S}_{\mathrm{rs}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{F}_{1}^{*}\times\mathbf{F}_{1}\right]=\frac{1}{\omega c}\mathbf{P}$, where $\mathbf{P}=\frac{c}{8\pi}\operatorname{Re}\left[\mathbf{E}^{*}\times\mathbf{H}\right]$ is the time-averaged Poynting vector. This RS spin characterizes the “chirality” of power flow, and its “handedness” is defined by $\kappa=\pm 1$.
We further consider a plane wave with circularly polarized electric and magnetic fields: $\bm{\Psi}=[\left(\hat{\mathbf{e}}_{x}+\mathrm{i}\sigma_{\mathrm{e}}\hat{\mathbf{e}}_{y}\right)+\mathrm{i}\kappa(\hat{\mathbf{h}}_{y}-\mathrm{i}\sigma_{\mathrm{m}}\hat{\mathbf{h}}_{x})]Ee^{\mathrm{i}\kappa k_{0}z-\mathrm{i}\omega t}$, where $\sigma_{\mathrm{e}}=\pm 1$ denotes the electric spin, $\sigma_{\mathrm{m}}=\pm 1$ denotes the magnetic spin, and $\kappa=\pm 1$ denotes the RS spin. Note that $\sigma_{\mathrm{e}}=\sigma_{\mathrm{m}}$ for plane waves. The total spin density can be expressed as $\mathbf{S}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\bm{\Psi}^{*}\times\bm{\Psi}\right]=\mathbf{S}_{\mathrm{e}}+\mathbf{S}_{\mathrm{m}}+\mathbf{S}_{\mathrm{rs}}$, where $\mathbf{S}_{\mathrm{e}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{E}^{*}\times\mathbf{E}\right]$ is the electric spin density, $\mathbf{S}_{\mathrm{m}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{H}^{*}\times\mathbf{H}\right]$ is the magnetic spin density, and $\mathbf{S}_{\mathrm{rs}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{F}_{1}^{*}\times\mathbf{F}_{1}\right]+\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{F}_{2}^{*}\times\mathbf{F}_{2}\right]=\frac{1}{\omega c}\mathbf{P}$ is the RS spin density. Notably, the sum of $\mathbf{S}_{\mathrm{e}}$ and $\mathbf{S}_{\mathrm{m}}$ corresponds to the conventional optical spin density that has been extensively studied in recent years [14, 49, 50, 51, 52, 53, 54], while the RS spin density $\mathbf{S}_{\mathrm{rs}}$ has been largely overlooked [48, 55]. Figure 1e-h shows the instantaneous $\mathbf{E}$, $\mathbf{H}$, $\mathbf{F}_{1}$, and spin directions for the plane waves with $\left(\sigma_{\mathrm{e}},\sigma_{\mathrm{m}},\kappa\right)=(+1,+1,+1),(+1,+1,-1),(-1,-1,+1)$, and (1,1,1)(-1,-1,-1), respectively. Their 4D polarizations are represented on the Poincaré hypersphere in Fig. 1a as the points “e”, “f”, “g”, and “h”, respectively.
The evolution of 4D polarization state $|\Psi\rangle$ simultaneously traces out paths on the E-sphere, H-sphere, and RS-sphere. The electric (magnetic) polarization evolution can generate the PB phase $\Phi_{\mathrm{pb}}$, which is proportional to the solid angle subtended by the area enclosed by the evolution path on E-sphere (H-sphere) [1]. Notably, the electric and magnetic polarization evolutions are intrinsically linked through the Maxwell equations. For the considered paraxial waves, the electric and magnetic polarization evolutions trace the same path and contribute to the same PB phase. Additionally, a new type of geometric phase can arise from the evolution of RS polarization (i.e., the polarization of $\mathbf{F}_{1,2}$). This geometric phase, termed RS phase and denoted as $\Phi_{\mathrm{rs}}$, is proportional to the solid angle subtended by the area enclosed by the evolution path on the RS-sphere (Supplementary Note II). Importantly, $\Phi_{\mathrm{rs}}$ is independent of $\Phi_{\mathrm{pb}}$ because $\Phi_{\mathrm{pb}}$ manifests in the basis states $\left|H_{\mathrm{rs}}\right\rangle$ and $\left|V_{\mathrm{rs}}\right\rangle$. Therefore, the total geometric phase due to 4D polarization evolution is $\Phi_{\mathrm{tot}}=\Phi_{\mathrm{pb}}+\Phi_{\mathrm{rs}}$.
[…]
RS geometric phase at interfaces

### v2

# Riemann-Silberstein geometric phase in 4D polarization space

@@ Abstract
Geometric phase is a far-reaching concept in quantum and classical physics. The first discovered geometric phase, the Pancharatnam-Berry (PB) phase, has profoundly shaped nanophotonics through metasurfaces. However, the PB phase arises from SU(2) polarization evolution and is constrained to a 2D polarization space, failing to capture the full polarization degrees of freedom. We generalize geometric phase to the 4D Riemann-Silberstein (RS) space that simultaneously describes electric, magnetic, and hybrid electric-magnetic polarizations. We show that SU(4) polarization evolution can generate a new geometric phase, the RS phase, alongside the PB phase. Unlike the PB phase that typically manifests in circularly polarized light, the RS phase can emerge in arbitrarily polarized light. Together, they enable a high-dimensional geometric framework for light propagation across general interfaces. We reveal that the phase shifts governed by Fresnel equations are direct manifestations of the RS-space geometric phases, integrating a century-old wave theory into this paradigm. We experimentally validate the framework using metasurfaces and achieve high-dimensional wavefront manipulation. Our work offers fundamental insights into the geometric nature of light-matter interactions, with implications for topological and non-Abelian physics in classical wave systems.
@@ Introduction
Geometric phases emerge from state evolution in parameter space and provide a unified framework for understanding diverse phenomena in quantum and classical physics [RN54, RN53, shapere1989geometric, RN51]. These phases have been extensively studied in various physical systems, including quantum particles [RN53, RN49, RN58, RN48], condensed matter [RN47, RN46], and classical wave systems [RN45, chiao1986manifestations, RN43, RN42, RN41]. In optics, geometric phases can give rise to intriguing phenomena such as spin-orbit interactions [RN40] and photonic topological states [RN39, RN38, RN37, RN36], providing important insights into the geometric and topological properties of optical fields [RN35, RN34, RN33] and enabling novel mechanisms for light manipulation [RN32, RN31, RN30, RN29, kim2022metasurface, xiong2023breaking].
The PB phase has recently attracted significant attention for its crucial role in topological meta-optics and structured light manipulation [RN27, RN26, RN25, RN24, RN23, song2021plasmonic, RN22, RN21, zeng2025metalasers]. This phase arises from SU(2) evolution of electric polarization on the Poincaré sphere [RN20, RN19], which is a 2D space describing the polarization of a two-component spinor. However, light is an electromagnetic wave comprising an electric field $\mathbf{E}$ and a magnetic field $\mathbf{H}$, which can be described as a bispinor in the Dirac-like formulations of Maxwell’s equations [barnett2014optical, alpeggiani2018]. For monochromatic waves, both $\mathbf{E}$ and $\mathbf{H}$ are two-component vector fields within the local frame defined by their polarization ellipses [RN35, RN18], and their polarizations can be different. Consequently, the complete polarization state of a monochromatic electromagnetic wave resides in a 4D Hilbert space (i.e., direct sum of electric and magnetic polarization spaces), termed RS space in this paper, which simultaneously characterizes electric, magnetic, and hybrid electric-magnetic polarizations. The SU(4) evolution of the 4D complete polarization can give rise to nontrivial geometric phases beyond the conventional PB phase, which have thus far remained elusive. These RS-space geometric phases are essential to establishing a unified geometric framework for light-matter interactions involving complex materials or structured light.
@@
In this work, we expand the geometric phase paradigm from the 2D polarization space to the 4D polarization space. We identify a new class of geometric phase, the RS phase, which can emerge in general electromagnetic waves and is governed by the hybrid electric-magnetic polarization, in contrast to the conventional PB phase associated with individual electric (magnetic) polarization. The RS and PB phases complement each other, and together enable a high-dimensional geometric framework for light propagation across general interfaces. This framework reveals a hidden geometric nature of the phase shifts due to light transmission and reflection at interfaces, a classical phenomenon governed by Fresnel equations. We experimentally demonstrate the RS phase and verify the framework at microwave frequencies using metasurfaces, which gives rise to high-dimensional wavefront deflection that cannot be achieved in conventional metasurface systems. The results provide new insights into light-matter interactions at interfaces and expand the toolkit for exploring geometric and topological properties of light.
Geometric phases in RS space
@@
The polarization state of light is conventionally described as a two-component spinor governed by an effective Hamiltonian $\hat{\mathcal{H}}=\hat{\rho}-\frac{1}{2}\hat{\text{I}}_{2}=\frac{1}{2}\mathbf{S}\cdot\hat{\bm{\sigma}}$, where $\hat{\rho}=|E\rangle\langle E|$ is the polarization projector, $\hat{\text{I}}_{2}$ is the $2\times 2$ identity matrix, $\hat{\bm{\sigma}}$ is the vector of Pauli spin matrices, and $\mathbf{S}$ is the Stokes vector [RN20]. The 2D polarization state $|E\rangle$ can be represented on the Poincaré sphere [RN54]. In our framework, a monochromatic electromagnetic field is described by the wavefuntion $\bm{\Psi}=\mathbf{E}+\mathrm{i}\mathbf{H}$ (Gaussian units are used throughout the paper), as in the Dirac-like formulation of electromagnetism [bialynicki1996v]. Here, $\mathbf{E}(\mathbf{r},t)=E_{i}(\mathbf{r},t)\hat{\mathbf{e}}_{i}+E_{j}(\mathbf{r},t)\hat{\mathbf{e}}_{j}$ is the complex electric field in the local frame with bases $\left(\hat{\mathbf{e}}_{i},\hat{\mathbf{e}}_{j}\right)$, and $\mathbf{H}(\mathbf{r},t)=H_{i}(\mathbf{r},t)\hat{\mathbf{h}}_{i}+H_{j}(\mathbf{r},t)\hat{\mathbf{h}}_{j}$ is the complex magnetic field in the local frame with bases $(\hat{\mathbf{h}}_{i},\hat{\mathbf{h}}_{j})$. The complete polarization state $|\Psi\rangle$ is a four-component bispinor governed by the effective Hamiltonian $\hat{\mathcal{H}}=\hat{\rho}-\frac{1}{4}\hat{\text{I}}_{4}=\frac{1}{2}\mathbf{S}\cdot\hat{\bm{\lambda}}$, where $\hat{\rho}=|\Psi\rangle\langle\Psi|$ and $\hat{\bm{\lambda}}$ is the generalized Gell-Mann matrices (Supplementary Note I). This 4D polarization state exhibits both internal polarizations (i.e., the polarizations of individual $\mathbf{E}$ and $\mathbf{H}$ fields) and external polarization (i.e., the hybrid $\mathbf{E}$-$\mathbf{H}$ polarization). It resides in a 4D RS Hilbert space and can be represented on a Poincaré hypersphere [RN17, RN16], as illustrated in Fig. 1a, which comprises three nested Poincaré spheres: E-sphere characterizing the $\mathbf{E}$ field polarization, H-sphere characterizing the $\mathbf{H}$ field polarization, and RS-sphere characterizing the hybrid $\mathbf{E}$-$\mathbf{H}$ polarization. The E-sphere has the north pole state $\left|N_{\mathrm{e}}\right\rangle$ corresponding to right-handed circularly-polarized (RCP) electric field and the south pole state $\left|S_{\mathrm{e}}\right\rangle$ corresponding to left-handed circularly-polarized (LCP) electric field; The H-sphere has the north pole state $\left|N_{\mathrm{m}}\right\rangle$ corresponding to RCP magnetic field and the south pole state $\left|S_{\mathrm{m}}\right\rangle$ corresponding to the LCP magnetic field. Any other point on the E-sphere or H-sphere denotes a $\mathrm{SU}(2)$ superposition state parameterized by the polar angle $\theta_{\mathrm{e},\mathrm{m}}$ and azimuthal angle $\phi_{\mathrm{e},\mathrm{m}}$. Notably, the RS-sphere characterizes the relationship between E-sphere and H-sphere with $\theta_{\mathrm{rs}}$ and $\phi_{\mathrm{rs}}$; its horizontal basis state $\left|H_{\mathrm{rs}}\right\rangle$ corresponds to an arbitrary state on the E-sphere; its vertical basis state $\left|V_{\mathrm{rs}}\right\rangle$ corresponds to an arbitrary state on the H-sphere. Any other point on the RS-sphere denotes a $\operatorname{SU}(4)$ superposition state $|\Psi\rangle$ (see Methods). The SU(4) evolution of $|\Psi\rangle$ will generally trace out trajectories on all the three spheres, and each sphere with an effective magnetic monopole in the center will contribute to the total geometric phase of $\bm{\Psi}$ with a weighting. In particular, the polarization evolution on the RS-sphere will generate a new type of geometric phase, termed RS phase, which will be elaborated with concrete examples below.
@@
In the following, we will focus on paraxial waves involving transverse electric and magnetic fields that are mutually perpendicular. In such cases, $\mathbf{E}$ and $\mathbf{H}$ exhibit the same polarization, and $\bm{\Psi}$ can be decomposed into a pair of orthogonal RS vectors: $\bm{\Psi}=\mathbf{F}_{1}+\mathbf{F}_{2}$, where $\mathbf{F}_{1}(\mathbf{r},t)=E_{i}(\mathbf{r},t)\hat{\mathbf{e}}_{i}+\mathrm{i}H_{j}(\mathbf{r},t)\hat{\mathbf{h}}_{j}$ and $\mathbf{F}_{2}(\mathbf{r},t)=E_{j}(\mathbf{r},t)\hat{\mathbf{e}}_{j}+\mathrm{i}H_{i}(\mathbf{r},t)\hat{\mathbf{h}}_{i}$. Consequently, the hybrid $\mathbf{E}$-$\mathbf{H}$ polarization reduces to the polarizations of $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$. We note that the RS vectors $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ are composed of complex electric and magnetic fields, which are different from the conventional RS vector comprising real electric and magnetic fields [RN15, RN13, RN57, RN12, RN11].
We use electromagnetic plane waves as an example to illustrate the 4D polarization and its representation on the Poincaré hypersphere. We first consider a plane wave with linearly polarized electric and magnetic fields: $\bm{\Psi}=(\hat{\mathbf{e}}_{x}+\mathrm{i}\kappa\hat{\mathbf{h}}_{y})Ee^{\mathrm{i}\kappa k_{0}z-\mathrm{i}\omega t}$, where $\kappa=\pm 1$ denotes the sign of the propagation direction relative to +z+z axis. Figures 1c and 1d show the instantaneous $\mathbf{E}$, $\mathbf{H}$, and $\mathbf{F}_{1}$ fields for $\kappa=1$ and $\kappa=-1$, respectively. The polarizations of $\mathbf{E}$ and $\mathbf{H}$ fields are represented by points “c” and “d” on the E-sphere and H-sphere in Fig. 1a. The RS field $\mathbf{F}_{1}$ is RCP for $\kappa=1$ and LCP for $\kappa=-1$, which are represented by points “c” and “d” on the RS-sphere in Fig. 1a. Naturally, we can introduce an RS spin density ${\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{rs}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{F}_{1}^{*}\times\mathbf{F}_{1}\right]=\frac{1}{\omega c}\mathbf{P}$, where $\mathbf{P}=\frac{c}{8\pi}\operatorname{Re}\left[\mathbf{E}^{*}\times\mathbf{H}\right]$ is the time-averaged Poynting vector. This RS spin originates from the intrinsic chirality of RS fields with the handedness defined by $\kappa=\pm 1$.
We further consider a plane wave with circularly polarized electric and magnetic fields: $\bm{\Psi}=[\left(\hat{\mathbf{e}}_{x}+\mathrm{i}\sigma_{\mathrm{e}}\hat{\mathbf{e}}_{y}\right)+\mathrm{i}\kappa(\hat{\mathbf{h}}_{y}-\mathrm{i}\sigma_{\mathrm{m}}\hat{\mathbf{h}}_{x})]Ee^{\mathrm{i}\kappa k_{0}z-\mathrm{i}\omega t}$, where $\sigma_{\mathrm{e}}=\pm 1$ denotes the electric spin, $\sigma_{\mathrm{m}}=\pm 1$ denotes the magnetic spin, and $\kappa=\pm 1$ denotes the RS spin. Note that $\sigma_{\mathrm{e}}=\sigma_{\mathrm{m}}$ for plane waves. The total spin density can be expressed as ${\color[rgb]{0,0,0}\mathbf{s}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\bm{\Psi}^{*}\times\bm{\Psi}\right]={\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{e}}+{\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{m}}+{\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{rs}}$, where ${\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{e}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{E}^{*}\times\mathbf{E}\right]$ is the electric spin density, ${\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{m}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{H}^{*}\times\mathbf{H}\right]$ is the magnetic spin density, and ${\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{rs}}=\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{F}_{1}^{*}\times\mathbf{F}_{1}\right]+\frac{1}{16\pi\omega}\operatorname{Im}\left[\mathbf{F}_{2}^{*}\times\mathbf{F}_{2}\right]=\frac{1}{\omega c}\mathbf{P}$ is the RS spin density. Notably, the sum of ${\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{e}}$ and ${\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{m}}$ corresponds to the conventional optical spin density that has been extensively studied in recent years [RN40, RN10, RN8, RN7, RN6, RN5, RN4], while the RS spin density ${\color[rgb]{0,0,0}\mathbf{s}}_{\mathrm{rs}}$ has been largely overlooked [RN11, golat2025electromagnetic]. Figure 1e-h shows the instantaneous $\mathbf{E}$, $\mathbf{H}$, $\mathbf{F}_{1}$, and spin directions for the plane waves with $\left(\sigma_{\mathrm{e}},\sigma_{\mathrm{m}},\kappa\right)=(+1,+1,+1),(+1,+1,-1),(-1,-1,+1)$, and (1,1,1)(-1,-1,-1), respectively. Their 4D polarizations are represented on the Poincaré hypersphere in Fig. 1a as the points “e”, “f”, “g”, and “h”, respectively.
The evolution of 4D polarization state $|\Psi\rangle$ simultaneously traces out paths on the E-sphere, H-sphere, and RS-sphere. The electric (magnetic) polarization evolution can generate the PB phase $\Phi_{\mathrm{pb}}$, which is proportional to the solid angle subtended by the area enclosed by the evolution path on E-sphere (H-sphere) [RN54]. Notably, the electric and magnetic polarization evolutions are intrinsically linked through the Maxwell equations. For the considered paraxial waves, the electric and magnetic polarization evolutions trace the same path and contribute to the same PB phase. Additionally, the polarization evolution of $\mathbf{F}_{1,2}$ will generate the RS phase $\Phi_{\mathrm{rs}}$, which is proportional to the solid angle subtended by the area enclosed by the evolution path on the RS-sphere (Supplementary Note III). Importantly, $\Phi_{\mathrm{rs}}$ is independent of $\Phi_{\mathrm{pb}}$ because $\Phi_{\mathrm{pb}}$ manifests in the basis states $\left|H_{\mathrm{rs}}\right\rangle$ and $\left|V_{\mathrm{rs}}\right\rangle$. Therefore, the total geometric phase due to 4D polarization evolution is $\Phi_{\mathrm{tot}}=\Phi_{\mathrm{pb}}+\Phi_{\mathrm{rs}}$.
[…]
@@
RS phase at interfaces